the length of a rectangle is 2 centimeters less than 3 times its width. its area is 21 square centimeters find the dimensions of the rectangle
Answer by mananth(16946)

the length of a rectangle is 2 centimeters less than 3 times its width. its area is 21 square centimeters find the dimensions of the rectangle
..
width =w
length = 3w-2
...
L*W= Area
w(3w-2)=21
3w^2-2w=21
3w^2-2w-21=0
3w^2-9w+7w-21=0
3w(w-3)+7(w-3)=0
(w-3)(3w+7)=0
..
w=3 cm the width
length = 3w-2
Length =7 cm
...
m.ananth@hotmail.ca